You are the manager of College Computers, a manufacturer of customized computers that meet the specifications required by the local university. Over 90 percent of your clientele consists of college students. College Computers is not the only firm that builds computers to meet this university’s specifications; indeed, it competes with many manufacturers online and through traditional retail outlets. To attract its large student clientele, College Computers runs a weekly ad in the student paper advertising its “free service after the sale” policy in an attempt to differentiate itself from the competition. The weekly demand for computers produced by College Computers is given by Q = 800 – 2P, and its weekly cost of producing computers is C(Q) = 1,200 + 2Q2.

If other firms in the industry sell PCs at $300, what price and quantity of computers should you produce to maximize your firm’s profits?

Price: $ 360 Correct
Quantity: 80 Correct computers

What long-run adjustments should you anticipate?

I need help figuring out how find MC? I know MC=4Q but I want to know how to find that? Thank you in advance!
Answer #1

Marginal cost is the derivative of the total cost with respect to the quantity. In this case the total cost function is C = 1200 + 2Q^2

Therefore we have MC = dMC / dQ = 2*2Q^(2-1)

= 4Q^1

= 4Q

This shows that marginal cost is 4Q.